OpenStudy (anonymous):

i need help !!!! please cos 2θ = 13 sin θ − 6 even if you just give me the answer that would be great

OpenStudy (anonymous):

Are you trying to prove the identity, or solve for tetha?

OpenStudy (anonymous):

solve for theta

OpenStudy (anonymous):

Is there a domain? ex, from 0 to 2pi?

OpenStudy (anonymous):

nope it just says solve the equation

OpenStudy (anonymous):

Very well, This shouldn't be hard then.

OpenStudy (anonymous):

So, start by moving everything to the left side. You get 6+cos(2x)-13sin(x)=0 (i'll use x instead of theta to make typing easier.)

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

now using the identity 1-sin^2(x), transform cos(2x). You get something like 7-13sin(x)-2sin^2x=0

OpenStudy (anonymous):

Now factor

OpenStudy (anonymous):

Should i continue?

OpenStudy (anonymous):

Let me know, otherwise there are many others that requested my help.

OpenStudy (anonymous):

yes please

OpenStudy (anonymous):

Alright, so when I factored I get: -((7+sin(x))(2sin(x)-1)=0 Multiply both sides by -1 to eliminate the negative.

OpenStudy (anonymous):

Now you have 2 equations 7+sinx=0 or 2sinx-1=0

OpenStudy (anonymous):

okay so sinx = 0 a

OpenStudy (anonymous):

so solve the first equation, sinx=-7 take the inverse sine of -7. This looks like a mess, so just see what the other equation produces.

OpenStudy (anonymous):

(2sin(x)-1)=0 add the 1, then divide by 2 sinx=(1/2) x=5pi/6, and x=pi/6

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

So your two answers are posted above. Hope that helped. If still confused just ask.

OpenStudy (anonymous):

but it would be theta = ........ and ............

OpenStudy (anonymous):

that was the answer thank you !

OpenStudy (anonymous):

sure thing.